# Sampling from the Mean-Field Stationary Distribution

### v2

@@ B.3 Logarithmic Sobolev Inequalities via Perturbations
$C_{\mathsf{LSI}}\bigl{(}\mu^{i|-i}(\cdot\mid x^{-i})\bigr{)}\leq\frac{4\sigma^{2}}{\lambda}+\Bigl{(}\frac{2B}{\lambda}+\frac{\sigma}{\sqrt{\lambda}}\Bigr{)}^{2}\,\Bigl{(}2+d+\frac{16B^{2}}{\lambda\sigma^{2}}\Bigr{)}\exp\Bigl{(}\frac{2B^{2}}{\lambda\sigma^{2}}\Bigr{)}\,.$
Applying the tensorization argument from the proof of Lemma 8 completes the proof. ∎
@@ Appendix C Explicit Calculations for the Gaussian Case
$\qquad{}-{(I_{N}\otimes A_{\lambda})}^{-1}(1_{N}\otimes I_{d})\bigl{(}I_{d}+(1_{N}^{\mathsf{T}}\otimes I_{d})\mathopen{}\mathclose{{}\left(I_{N}\otimes A_{\lambda}}\right)^{-1}(1_{N}\otimes I_{d})\bigr{)}^{-1}(1_{N}^{\mathsf{T}}\otimes I_{d})\mathopen{}\mathclose{{}\left(I_{N}\otimes A_{\lambda}}\right)^{-1}$
@@ Appendix D Additional Technical Lemmas
Lemma 24 (Information Inequality [Csi84]).
@@ Appendix E Sampling Guarantees
Lemma 25 ([Zha+23, Proof of Theorem 6]).
@@
  • LMC: We use the guarantee for Langevin Monte Carlo from [VW19].
  • MALA–PS: We use the guarantee for the Metropolis-adjusted Langevin algorithm together with the proximal sampler from [AC23]. Note that the iteration complexity is $\widetilde{\mathcal{O}}(\kappa d^{1/2}N^{1/2})$, and we substitute in the chosen value for $N$.
  • ULMC–PS: Here, we use underdamped Langevin Monte Carlo to implement the proximal sampler. To justify the sampling guarantee, note that since $\log\mu^{1:N}$ is $\beta$-smooth, if we choose step size $h=\frac{1}{2\beta}$ for the proximal sampler, then the RGO is $\beta$-strongly log-concave and $3\beta$-log-smooth. According to [AC23, Proof of Theorem 5.3], it suffices to implement the RGO in each iteration to accuracy $N^{1/2}\varepsilon/\kappa^{1/2}$ in $\sqrt{\operatorname{\mathsf{KL}}}$. Then, from [Zha+23], this can be done via ULMC with complexity $\widetilde{\mathcal{O}}(\kappa^{1/2}d^{1/2}/\varepsilon)$. Finally, since the number of outer iterations of the proximal sampler is $\widetilde{\mathcal{O}}(\kappa)$, we obtain the claim.
  • ${}^{+}$: Here, we use either the randomized midpoint discretization [SL19] or the shifted ODE discretization [FLO21] of the underdamped Langevin diffusion. We also replace the LSI assumptions (Assumptions 2 and 9) with strong convexity (Assumption 4).
@@
  • LMC: For Langevin Monte Carlo, we use the $\chi^{2}$ guarantee from [Che+22a] and the $\mathcal{W}_{2}$ guarantee from [DMM19].
  • ULMC: For underdamped Langevin Monte Carlo, we use the $\chi^{2}$ guarantee from [AC23].
  • ${}^{+}$: Here, we use the $\mathcal{W}_{2}$ guarantees for either the randomized midpoint discretization [SL19] or the shifted ODE discretization [FLO21] of the underdamped Langevin diffusion.

### v4

@@ B.3 Logarithmic Sobolev Inequalities via Perturbations
$\qquad{}\mapsto\sum_{i=1}^{N}\int\Bigl{[}\frac{\sigma^{2}\,\lVert y^{i}-x^{i}\rVert^{2}}{4t}-\bigl{(}\frac{\lambda}{2}-\frac{\sigma^{2}}{4}\bigr{)}\,\lVert x^{i}\rVert^{2}\Bigr{]}\,\pi^{i}(\mathrm{d}x^{i})+N\mathcal{F}_{0}\bigl{(}\frac{1}{N}\sum_{i=1}^{N}\pi^{i}\bigr{)}+\frac{\sigma^{2}}{2}\sum_{i=1}^{N}\operatorname{\mathsf{ent}}(\pi^{i})$
satisfies the system of equations (B.3), and that the minimizer is unique because the functional is strictly convex. We also let $\pi_{t,y^{1:N}}\coloneqq\bigotimes_{i=1}^{N}\pi_{t,y^{i}}$.
Then, for any unit vector $\theta^{1:N}\in\mathbb{R}^{d\times N}$,
$\langle\theta^{1:N},\operatorname{cov}_{\mu_{t,y^{1:N}}}\theta^{1:N}\rangle$ $\leq\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu_{t,y^{1:N}}}[\langle\theta^{1:N},x^{1:N}-\operatorname{\mathbb{E}}_{\bar{x}^{1:N}\sim\pi_{t,y^{1:N}}}\bar{x}^{1:N}\rangle^{2}]$
$\leq\Bigl{(}\mathcal{W}_{2}(\mu_{t,y^{1:N}},\pi_{t,y^{1:N}})+\sqrt{\langle\theta^{1:N},\operatorname{cov}_{\pi_{t,y^{1:N}}}\theta^{1:N}\rangle}\Bigr{)}^{2}$
$\leq\Bigl{(}\mathcal{W}_{2}(\mu_{t,y^{1:N}},\pi_{t,y^{1:N}})+\max_{i\in[N]}\sqrt{\lVert\operatorname{cov}_{\pi_{t,y^{i}}}\rVert_{\mathsf{op}}}\Bigr{)}^{2}\,,$
where we used the fact that $\pi_{t,y^{1:N}}$ is a product measure. Also, introduce
$\breve{\pi}_{t,y^{i}}(x^{i})\propto\exp\Bigl{(}-\frac{\lVert y^{i}-x^{i}\rVert^{2}}{2t}-\bigl{(}\frac{\lambda}{\sigma^{2}}-\frac{1}{2}\bigr{)}\,\lVert x^{i}\rVert^{2}\Bigr{)}\,,$
so that $\pi_{t,y^{i}}\propto\exp(-\frac{2}{\sigma^{2}}\,\delta\mathcal{F}_{0}(\bar{\pi}_{t,y^{1:N}},\cdot))\,\breve{\pi}_{t,y^{i}}$. The same argument as above then yields
$\langle\theta^{i},\operatorname{cov}_{\pi_{t,y^{i}}}\theta^{i}\rangle\leq\Bigl{(}\mathcal{W}_{2}(\pi_{t,y^{i}},\breve{\pi}_{t,y^{i}})+\sqrt{\langle\theta^{i},\operatorname{cov}_{\breve{\pi}_{t,y^{i}}}\theta^{i}\rangle}\Bigr{)}^{2}\,.$
Since $\breve{\pi}_{t,y^{i}}$ is $(2\lambda/\sigma^{2}-1+1/t)$-strongly log-concave, $\lVert\operatorname{cov}_{\breve{\pi}_{t,y^{i}}}\rVert_{\mathsf{op}}\leq 1/(2\lambda/\sigma^{2}-1+1/t)$ by the Brascamp–Lieb inequality [BraLie1976]. Also, since $\frac{2}{\sigma^{2}}\,\delta\mathcal{F}_{0}(\bar{\pi}_{t,y^{1:N}},\cdot)$ is $2B/\sigma^{2}$-Lipschitz under Assumption 8, then [KhuMaaPed24LInfOT, Corollary 2.4] yields
$\mathcal{W}_{2}(\pi_{t,y^{i}},\breve{\pi}_{t,y^{i}})$ $\leq\mathcal{W}_{\infty}(\pi_{t,y^{i}},\breve{\pi}_{t,y^{i}})\leq\frac{2B/\sigma^{2}}{2\lambda/\sigma^{2}-1+1/t}\,.$
Hence,
$\lVert\operatorname{cov}_{\pi_{t,y^{i}}}\rVert_{\mathsf{op}}$ $\leq\Bigl{(}\frac{1}{\sqrt{2\lambda/\sigma^{2}-1+1/t}}+\frac{2B/\sigma^{2}}{2\lambda/\sigma^{2}-1+1/t}\Bigr{)}^{2}\,.$
Finally, it remains to control $\mathcal{W}_{2}(\mu_{t,y^{1:N}},\pi_{t,y^{1:N}})$. Note that when $t=\infty$, this essentially reduces to Theorem 5, so the task is to prove a generalization thereof. Note that by Lemma 21, in Theorem 26 below, we can take
$\overline{C}_{\mathsf{LSI}}$ $\leq\frac{1}{2\lambda/\sigma^{2}-1+1/t}\exp\Bigl{(}\frac{4B^{2}/\sigma^{4}}{2\lambda/\sigma^{2}-1+1/t}+\frac{8B/\sigma^{2}}{\sqrt{2\lambda/\sigma^{2}-1+1/t}}\Bigr{)}$
$\leq\frac{3}{2\lambda/\sigma^{2}-1+1/t}\exp\Bigl{(}\frac{8B^{2}/\sigma^{4}}{2\lambda/\sigma^{2}-1+1/t}\Bigr{)}\leq\frac{3}{2\lambda/\sigma^{2}-1}\exp\Bigl{(}\frac{8B^{2}/\sigma^{4}}{2\lambda/\sigma^{2}-1}\Bigr{)}\,.$
In particular, under the assumption (B.2) for $N$, the preconditions of Theorem 26 are met and it yields the bound
$\mathcal{W}_{2}(\mu_{t,y^{1:N}},\pi_{t,y^{1:N}})$ $\lesssim\frac{\overline{C}_{\mathsf{LSI}}\sqrt{\beta d}}{\sigma}\lesssim\frac{\sqrt{\beta d}}{\sigma\,(2\lambda/\sigma^{2}-1+1/t)}\exp\Bigl{(}\frac{8B^{2}/\sigma^{4}}{2\lambda/\sigma^{2}-1}\Bigr{)}\,.$
Putting everything together completes the proof. ∎
Corollary 25 (Uniform LSI for the Stationary Measure).
Under Assumptions 56, and 8, if
$N\gtrsim\frac{\beta d}{\lambda}\exp\Bigl{(}\frac{8B^{2}}{\lambda\sigma^{2}}\Bigr{)}$ (B.4)
for a sufficiently large implied (but universal) constant, then
$C_{\mathsf{LSI}}(\mu^{1:N})$ $\lesssim\frac{\sigma^{2}}{\lambda}\exp\Bigl{(}\mathcal{O}\Bigl{(}\frac{B^{2}}{\lambda\sigma^{2}}+\frac{\beta d}{\lambda}\exp\frac{16B^{2}}{\lambda\sigma^{2}}\Bigr{)}\Bigr{)}\,.$
Proof.  To meet the conditions of Lemma 24, we perform a rescaling. We abuse notation and denote by $\eta:\mathbb{R}^{d}\to\mathbb{R}^{d}$ the scaling map $x\mapsto\eta x$. Then,
$\mu_{\eta}^{1:N}(x^{1:N})$ $\coloneqq\eta_{\#}\mu^{1:N}(x^{1:N})\propto\mu^{1:N}(\eta^{-1}x^{1:N})\propto\exp\Bigl{(}-\frac{2N}{\sigma^{2}}\,\mathcal{F}_{0}^{\eta}(\rho_{x^{1:N}})-\frac{\lambda}{\eta^{2}\sigma^{2}}\,\lVert x^{1:N}\rVert^{2}\Bigr{)}\,,$
where $\mathcal{F}_{0}^{\eta}(\nu)\coloneqq\mathcal{F}_{0}((\eta^{-1})_{\#}\nu)$. We see that $\mu_{\eta}^{1:N}$ is also the stationary measure for mean-field Langevin dynamics, with the following new parameters: $\beta\leftarrow\beta/\eta^{2}$; $\lambda\leftarrow\lambda/\eta^{2}$; $B\leftarrow B/\eta$. In particular, if we take $\eta^{2}=\lambda/\sigma^{2}$, then Lemma 24 applies to $\mu_{\eta}^{1:N}$ provided that (B.4) holds. Together with Lemma 23 with $a=1$ and $C\lesssim B/(\lambda^{1/2}\sigma)+{(\beta d/\lambda)}^{1/2}\exp(8B^{2}/(\lambda\sigma^{2}))$, it implies
$C_{\mathsf{LSI}}(\mu^{1:N}_{\eta})$ $\lesssim\exp\bigl{(}\mathcal{O}(C^{2})\bigr{)}=\exp\Bigl{(}\mathcal{O}\Bigl{(}\frac{B^{2}}{\lambda\sigma^{2}}+\frac{\beta d}{\lambda}\exp\frac{16B^{2}}{\lambda\sigma^{2}}\Bigr{)}\Bigr{)}\,.$
The result for $\mu^{1:N}$ follows from contraction mapping [bakry2014analysis, Proposition 5.4.3]. ∎
It remains to prove the following generalized propagation of chaos result.
Theorem 26 (Generalized Propagation of Chaos).
For each $i\in[N]$, let $V_{i}:\mathbb{R}^{d}\to\mathbb{R}$ and let $\mathcal{F}_{0}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$. Define probability measures
$\mu^{1:N}(x^{1:N})$ $\propto\exp\Bigl{(}-\frac{2}{\sigma^{2}}\sum_{i=1}^{N}V_{i}(x^{i})-\frac{2N}{\sigma^{2}}\,\mathcal{F}_{0}(\rho_{x^{1:N}})\Bigr{)}\,,$
$\pi^{i}(x^{i})$ $\propto\exp\Bigl{(}-\frac{2}{\sigma^{2}}\,V_{i}(x^{i})-\frac{2}{\sigma^{2}}\,\delta\mathcal{F}_{0}(\bar{\pi},x^{i})\Bigr{)}\,,$
where $\bar{\pi}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\pi_{i}$, $\pi^{1:N}\coloneqq\bigotimes_{i=1}^{N}\pi^{i}$, and we assume that these measures are all well-defined. Furthermore, assume that $\mathcal{F}_{0}$ satisfies Assumptions 5 and 6, and that for all $i\in[N]$ and all $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the proximal Gibbs measure
$\pi_{\nu}^{i}\propto\exp\Bigl{(}-\frac{2}{\sigma^{2}}\,\bigl{(}V_{i}+\delta\mathcal{F}_{0}(\nu,\cdot)\bigr{)}\Bigr{)}$
satisfies (LSI) with constant $\overline{C}_{\mathsf{LSI}}$. Then, for all $N\gtrsim\beta\overline{C}_{\mathsf{LSI}}d/\sigma^{2}$,
$\frac{1}{2\overline{C}_{\mathsf{LSI}}}\,\mathcal{W}_{2}^{2}(\mu^{1:N},\pi^{1:N})\leq\operatorname{\mathsf{KL}}(\mu^{1:N}\mathbin{\|}\pi^{1:N})\lesssim\frac{\beta\overline{C}_{\mathsf{LSI}}d}{\sigma^{2}}\,.$
Proof.  Since the proof closely follows the proof of Theorem 5, to avoid repetition we only highlight the main changes. We define the following energy functionals:
$\mathcal{E}^{N}(\nu^{1:N})$ $\coloneqq\sum_{i=1}^{N}\int V_{i}\,\mathrm{d}\nu^{i}+N\int\mathcal{F}_{0}(\rho_{x^{1:N}})\,\nu^{1:N}(\mathrm{d}x^{1:N})+\frac{\sigma^{2}}{2}\operatorname{\mathsf{ent}}(\nu^{1:N})\,,$
$\mathcal{E}(\nu^{1:N})$ $\coloneqq\sum_{i=1}^{N}\int V_{i}\,\mathrm{d}\nu^{i}+N\mathcal{F}_{0}(\bar{\nu})+\frac{\sigma^{2}}{2}\operatorname{\mathsf{ent}}(\nu^{1:N})\,,$
where for a measure $\nu^{1:N}$ on $\mathbb{R}^{d\times N}$, we use the notation $\bar{\nu}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\nu^{i}$ for the average of the marginals. The first step is to establish the analogue of the entropy toast inequality (Lemma 19). Letting $Z\coloneqq\prod_{i=1}^{N}\int\exp(-\frac{2}{\sigma^{2}}\,V_{i}-\frac{2}{\sigma^{2}}\,\delta\mathcal{F}_{0}(\bar{\pi},\cdot))$ denote the normalizing constant for $\pi^{1:N}$,
$\operatorname{\mathsf{KL}}(\nu^{1:N}\mathbin{\|}\pi^{1:N})$ $=\frac{2}{\sigma^{2}}\sum_{i=1}^{N}\int V_{i}\,\mathrm{d}\nu^{i}+\frac{2N}{\sigma^{2}}\iint\delta\mathcal{F}_{0}(\bar{\pi},z)\,\rho_{x^{1:N}}(\mathrm{d}z)\,\nu^{1:N}(\mathrm{d}x^{1:N})+\operatorname{\mathsf{ent}}(\nu^{1:N})$
$\qquad{}+\log Z$
$=\frac{2}{\sigma^{2}}\sum_{i=1}^{N}\int V_{i}\,\mathrm{d}(\nu^{i}-\pi^{i})+\frac{2N}{\sigma^{2}}\operatorname{\mathbb{E}}_{x^{1:N}\sim\nu^{1:N}}\int\delta\mathcal{F}_{0}(\bar{\pi},\cdot)\,\mathrm{d}(\rho_{x^{1:N}}-\bar{\pi})$
$\qquad{}+\operatorname{\mathsf{ent}}(\nu^{1:N})-\operatorname{\mathsf{ent}}(\pi^{1:N})$
$\leq\frac{2}{\sigma^{2}}\sum_{i=1}^{N}\int V_{i}\,\mathrm{d}(\nu^{i}-\pi^{i})+\frac{2N}{\sigma^{2}}\operatorname{\mathbb{E}}_{x^{1:N}\sim\nu^{1:N}}[\mathcal{F}_{0}(\rho_{x^{1:N}})-\mathcal{F}_{0}(\bar{\pi})]$
$\qquad{}+\operatorname{\mathsf{ent}}(\nu^{1:N})-\operatorname{\mathsf{ent}}(\pi^{1:N})$
$\leq\frac{2}{\sigma^{2}}\,\bigl{(}\mathcal{E}^{N}(\nu^{1:N})-\mathcal{E}(\pi^{1:N})\bigr{)}\,.$
From here, we find that
$\operatorname{\mathbb{E}}_{x^{1:N}\sim\nu^{1:N}}[\mathcal{F}_{0}(\rho_{x^{1:N}})-\mathcal{F}_{0}(\bar{\pi})]\leq\operatorname{\mathbb{E}}_{x^{1:N}\sim\nu^{1:N}}\int\delta\mathcal{F}_{0}(\rho_{x^{1:N}},\cdot)\,\mathrm{d}(\rho_{x^{1:N}}-\bar{\pi})\,.$
Now, moving on to the propagation of chaos part of this argument, we have
$\mathcal{E}^{N}(\mu^{1:N})-\mathcal{E}(\pi^{1:N})$
$\qquad\leq\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\Bigl{[}\int V_{i}\,\mathrm{d}(\mu^{i\mid-i}(\cdot\mid x^{-i})-\pi^{i})+\int\delta\mathcal{F}_{0}(\rho_{x^{1:N}},\cdot)\,\mathrm{d}(\rho_{x^{1:N}}-\pi^{i})\Bigr{]}$
$\qquad\qquad{}+\frac{\sigma^{2}}{2}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{-i}\sim\mu^{-i}}\bigl{[}\operatorname{\mathsf{ent}}(\mu^{i\mid-i}(\cdot\mid x^{-i}))-\operatorname{\mathsf{ent}}(\pi^{i})\bigr{]}$
$\qquad=\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\Bigl{[}\int V_{i}\,\mathrm{d}(\mu^{i\mid-i}(\cdot\mid x^{-i})-\pi^{i})+\delta\mathcal{F}_{0}(\rho_{x^{1:N}},x^{i})-\int\delta\mathcal{F}_{0}(\rho_{x^{1:N}},\cdot)\,\mathrm{d}\pi^{i}\Bigr{]}$
$\qquad\qquad{}+\frac{\sigma^{2}}{2}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{-i}\sim\mu^{-i}}\bigl{[}\operatorname{\mathsf{ent}}(\mu^{i\mid-i}(\cdot\mid x^{-i}))-\operatorname{\mathsf{ent}}(\pi^{i})\bigr{]}\,.$
To decouple, introduce a new variable $z^{i}\sim\mu^{i}$ independent of all the others and define as a shorthand $\tilde{x}_{i}^{1:N}$ as the vector $x^{1},\ldots,x^{i-1},z^{i},x^{i+1},\ldots,x^{N}$.
$\mathcal{E}^{N}(\mu^{1:N})-\mathcal{E}(\pi^{1:N})$
$\qquad\leq\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\Bigl{[}\int V_{i}\,\mathrm{d}(\mu^{i\mid-i}(\cdot\mid x^{-i})-\pi^{i})+\frac{\sigma^{2}}{2}\,\bigl{(}\operatorname{\mathsf{ent}}(\mu^{i\mid-i}(\cdot\mid x^{-i}))-\operatorname{\mathsf{ent}}(\pi^{i})\bigr{)}\Bigr{]}$
$\qquad\qquad{}+\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\Bigl{[}\operatorname{\mathbb{E}}_{z^{i}\sim\mu^{i}}\delta\mathcal{F}_{0}(\rho_{\tilde{x}^{1:N}_{i}},x^{i})-\int\operatorname{\mathbb{E}}_{z^{i}\sim\mu^{i}}\delta\mathcal{F}_{0}(\rho_{\tilde{x}^{1:N}_{i}},\cdot)\,\mathrm{d}\pi^{i}\Bigr{]}$
$\qquad{}+\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\Bigl{[}\delta\mathcal{F}_{0}(\rho_{x^{1:N}},x^{i})-\operatorname{\mathbb{E}}_{z^{i}\sim\mu^{i}}\delta\mathcal{F}_{0}(\rho_{\tilde{x}^{1:N}_{i}},x^{i})$
$\qquad\qquad\qquad\qquad\qquad{}-\int\bigl{(}\delta\mathcal{F}_{0}(\rho_{x^{1:N}},\cdot)-\operatorname{\mathbb{E}}_{z^{i}\sim\mu^{i}}\delta\mathcal{F}_{0}(\rho_{\tilde{x}^{1:N}_{i}},\cdot)\bigr{)}\,\mathrm{d}\pi^{i}\Bigr{]}\,.$
We group the terms in the first two lines as $\mathsf{A}$, and the terms in the last two lines as $\mathsf{B}$.
Let us first look at $\mathsf{A}$. If we introduce the proximal Gibbs measure for $\rho_{\tilde{x}^{1:N}_{i}}$ (in the $i$-th coordinate), defined via
$\tau_{\tilde{x}^{1:N}_{i}}^{i}\propto\exp\Bigl{(}-\frac{2}{\sigma^{2}}\,\bigl{(}V_{i}+\delta\mathcal{F}_{0}(\rho_{\tilde{x}_{i}^{1:N}},\cdot)\bigr{)}\Bigr{)}\,,$
one obtains as before
$\mathsf{A}\leq\frac{\sigma^{2}}{2}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\operatorname{\mathbb{E}}_{z^{i}\sim\mu^{i}}\operatorname{\mathsf{KL}}\bigl{(}\mu^{i\mid-i}(\cdot\mid x^{-i})\bigm{\|}\tau_{\tilde{x}^{1:N}_{i}}^{i}\bigr{)}\,.$
Applying the log-Sobolev inequality, it yields
$\mathsf{A}$ $\leq\frac{\beta^{2}\overline{C}_{\mathsf{LSI}}}{\sigma^{2}}\sum_{i=1}^{N}\operatorname{\mathbb{E}}_{x^{1:N}\sim\mu^{1:N}}\operatorname{\mathbb{E}}_{z^{i}\sim\mu^{i}}\mathcal{W}_{1}^{2}(\rho_{x^{1:N}},\rho_{\tilde{x}_{i}^{1:N}})\,.$
The Wasserstein distance is bounded by
$\mathcal{W}_{1}(\rho_{x^{1:N}},\rho_{\tilde{x}_{i}^{1:N}})$ $\leq\frac{1}{N}\,{\lVert x^{i}-z^{i}\rVert}\,.$
It eventually yields, as before,
$\mathsf{A}$ $\lesssim\frac{\beta^{2}\overline{C}_{\mathsf{LSI}}}{\sigma^{2}N}\,\Bigl{(}\frac{\overline{C}_{\mathsf{LSI}}}{N}\operatorname{\mathsf{KL}}(\mu^{1:N}\mathbin{\|}\pi^{1:N})+d\overline{C}_{\mathsf{LSI}}\Bigr{)}\,.$
As for the term $\mathsf{B}$, a straightforward modification of the proof of Theorem 5 readily yields
$\mathsf{B}\lesssim\frac{\beta\overline{C}_{\mathsf{LSI}}}{N}\operatorname{\mathsf{KL}}(\mu^{1:N}\mathbin{\|}\pi^{1:N})+\beta\overline{C}_{\mathsf{LSI}}d\,.$
Putting everything together yields the result. ∎
@@ Appendix C Explicit Calculations for the Gaussian Case
$\qquad{}-{(I_{N}\otimes A_{\lambda})}^{-1}(1_{N}\otimes I_{d})$
×$\qquad\qquad{}\times\bigl{(}I_{d}+(1_{N}^{\mathsf{T}}\otimes I_{d})\mathopen{}\mathclose{{}\left(I_{N}\otimes A_{\lambda}}\right)^{-1}(1_{N}\otimes I_{d})\bigr{)}^{-1}(1_{N}^{\mathsf{T}}\otimes I_{d})\mathopen{}\mathclose{{}\left(I_{N}\otimes A_{\lambda}}\right)^{-1}$
@@ Appendix D Additional Technical Lemmas
Lemma 28 (Information Inequality [csiszar1984sanov]).
@@ Appendix E Sampling Guarantees
Lemma 29 ([zhang2023improved, Proof of Theorem 6]).
@@
  • LMC: We use the guarantee for Langevin Monte Carlo from [vempala2019rapid].
  • MALA–PS: We use the guarantee for the Metropolis-adjusted Langevin algorithm together with the proximal sampler from [altschuler2023faster]. Note that the iteration complexity is $\widetilde{\mathcal{O}}(\kappa d^{1/2}N^{1/2})$, and we substitute in the chosen value for $N$.
  • ULMC–PS: Here, we use underdamped Langevin Monte Carlo to implement the proximal sampler. To justify the sampling guarantee, note that since $\log\mu^{1:N}$ is $\beta$-smooth, if we choose step size $h=\frac{1}{2\beta}$ for the proximal sampler, then the RGO is $\beta$-strongly log-concave and $3\beta$-log-smooth. According to [altschuler2023faster, Proof of Theorem 5.3], it suffices to implement the RGO in each iteration to accuracy $N^{1/2}\varepsilon/\kappa^{1/2}$ in $\sqrt{\operatorname{\mathsf{KL}}}$. Then, from [zhang2023improved], this can be done via ULMC with complexity $\widetilde{\mathcal{O}}(\kappa^{1/2}d^{1/2}/\varepsilon)$. Finally, since the number of outer iterations of the proximal sampler is $\widetilde{\mathcal{O}}(\kappa)$, we obtain the claim.
  • ULMC+: Here, we use either the randomized midpoint discretization [shen2019randomized] or the shifted ODE discretization [FosLyoObe21ShiftedODE] of the underdamped Langevin diffusion. We also replace the LSI assumptions (Assumptions 2 and 9) with strong convexity (Assumption 4).
@@
  • LMC: For Langevin Monte Carlo, we use the $\chi^{2}$ guarantee from [chewi2021analysis] and the $\mathcal{W}_{2}$ guarantee from [DurMajMia19LMCCvxOpt].
  • ULMC: For underdamped Langevin Monte Carlo, we use the $\chi^{2}$ guarantee from [altschuler2023faster].
  • ULMC+: Here, we use the $\mathcal{W}_{2}$ guarantees for either the randomized midpoint discretization [shen2019randomized] or the shifted ODE discretization [FosLyoObe21ShiftedODE] of the underdamped Langevin diffusion.